# Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory

### v1

@@ Abstract
The evaluation of large language models (LLMs) via benchmarks is widespread, yet inconsistencies between different leaderboards and poor separability among top models raise concerns about their ability to accurately reflect authentic model capabilities. This paper provides a critical analysis of benchmark effectiveness, examining main-stream prominent LLM benchmarks using results from diverse models. We first propose a new framework for accurate and reliable estimations of item characteristics and model abilities. Specifically, we propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture. Based on PSN-IRT, we conduct extensive analysis which reveals significant and varied shortcomings in the measurement quality of current benchmarks. Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference [^1] Code and data are openly available at https://github.com/Joe-Hall-Lee/PSN-IRT.
Lost in Benchmarks? Rethinking Large Language Model Benchmarking with Item Response Theory
[…]
footnotetext: * The first two authors contributed equally.
@@ 1 Introduction
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important. Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions aimed at assessing model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting deeper consideration of their effectiveness McIntosh et al. (2024).
As illustrated in Figure 1, on one hand, different benchmarks often produce inconsistent leaderboards Perlitz et al. (2024), with the same model achieving substantially varying rankings across evaluations. This makes it challenging for researchers and developers to determine a model’s real performance. On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^4] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets and its assumption of normally distributed abilities does not always hold.
@@
To validate the effectiveness of PSN-IRT, we conduct extensive experiments on evaluation results from 12 LLMs across 11 benchmarks. Our results show that PSN-IRT outperforms previous approaches in both parameter estimation accuracy and reliability. Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with key findings as follows:
@@
    We propose PSN-IRT, a benchmark analysis framework demonstrating superior estimation accuracy and reliability.
@@
    We conduct extensive experiments, using PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in may aspects.
@@ 2.1 Related Work
Researchers have explored various methods to enhance the reliability and interpretability of LLM benchmark evaluations. During test set construction, many benchmarks incorporate human annotation and leverage strong LLMs to ensure item quality, detect inconsistencies, and manage item difficulty He et al. (2024); Wei et al. (2024). Beyond these construction-phase strategies, recent studies have also introduced quantitative metrics to assess benchmark behavior post-deployment. For example, Benchmark Agreement Testing (BAT) evaluates the reliability of new benchmarks by comparing them with established ones using correlation-based measures Perlitz et al. (2024); White et al. (2025). Other work proposes novel metrics like Separability with Confidence for more nuanced benchmark analysis Li et al. (2024).
Item Response Theory (IRT) Lord et al. (1968); Baker and Kim (2004) offers a psychometric framework for finer-grained, item-level analysis. Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability Byrd and Srivastava (2022); Vania et al. (2021) and, more recently, to inspire methods for improving evaluation efficiency, for instance, through optimized item selection or adaptive testing paradigms Maia Polo et al. (2024); Zhuang et al. (2024). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored.
@@ 2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) Lord et al. (1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) DeVellis (2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
@@
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data Tsutsumi et al. (2021).
@@ 3 Pseudo-Siamese Network for IRT
  • The discriminability $a$ controls the steepness.
@@
4 Experiments
4.1 Set-up
Dataset.
@@ 4.1 Set-up
To construct training data, we conduct evaluations using OpenCompass Contributors (2023). The evaluation results are collected in the form of binary outcomes, indicating whether each model answered each item correctly. Then, we construct a unified outcome matrix by aggregating item-level results across all models and benchmarks. We split the interactions into training (60%), validation (20%), and test sets (20%). The validation set is utilized to monitor performance during model training and to implement early stopping, thereby preventing overfitting.
@@
ARC-C Clark et al. (2018) 295 General EM Multiple Choice
BBH Suzgun et al. (2023) 6,511 General EM Mixed
Chinese SimpleQA He et al. (2024) 3,000 Knowledge LLM-as-a-Judge QA
GPQA Diamond Rein et al. (2024) 198 Science EM Multiple Choice
GSM8K Cobbe et al. (2021) 1,319 Math EM QA
HellaSwag Zellers et al. (2019) 10,042 General EM Multiple Choice
HumanEval Chen et al. (2021) 164 Code Pass@1 QA
MATH Hendrycks et al. (2021b) 5,000 Math EM QA
MBPP Austin et al. (2021) 500 Code Pass@1 QA
MMLU Hendrycks et al. (2021a) 14,042 General EM Multiple Choice
TheoremQA Chen et al. (2023) 800 Science EM QA
@@
Model Parameter Method ACC F1 AUC Kendall Average
IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697 0.7989
MCMC 0.7070 0.7811 0.7278 0.9697 0.7964
VI 0.7201 0.8015 0.6940 0.9091 0.7812
VIBO 0.7188 0.8007 0.7055 0.9697 0.7987
Deep-IRT 1PL Deep Learning 0.7969 0.8507 0.8507 0.9697 0.8670
PSN-IRT 4PL Deep Learning 0.7991 0.8520 0.8477 1.0000 0.8747
@@
Model.
We mainly focus on currently high-performing models, including 360GPT2-Pro [^5] https://ai.360.cn/, DeepSeek-V3 DeepSeek-AI et al. (2025), Doubao-Pro [^6] https://www.volcengine.com/product/doubao, Gemini-1.5 Team et al. (2024a), Hunyuan-Turbo [^7] https://hunyuan.tencent.com/, Moonshot-v1888https://platform.moonshot.cn/, Qwen-Plus Yang et al. (2024), and Yi-Lightning Wake et al. (2025). However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality Martínez-Plumed et al. (2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it Team et al. (2024b), Mistral-7B-Instruct-v0.1 Jiang et al. (2023), Qwen2.5-3B-Instruct Yang et al. (2024), and Vicuna-7B-v1.3 Chiang et al. (2023).
Baseline.
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) Hastings (1970), Variational Inference (VI) Jordan et al. (1999), and VIBO Wu et al. (2020). We also compare with Deep-IRT Tsutsumi et al. (2021), which is a neural extension of 1PL IRT that learns item-trait interactions via deep networks.
Metric.
Two quantitative metrics are included:
  • Prediction accuracy: We measure how well each method predicts whether a model answers an item correctly. These metrics include accuracy, F1 score, and ROC AUC.
  • Rank stability: We assess the consistency of model rankings by splitting the test set into two subsets. For each subset, we estimate model abilities and generate a ranking. Kendall’s $\tau$ is then computed between the two rankings.
4.2 Estimation Ability of PSN-IRT
@@
Given our primary objective of analyzing test sets, all methods except Deep-IRT utilized the more comprehensive 4PL structure for modeling benchmark items. However, the experimental results indicate that employing different parameter structures has a minimal impact on prediction accuracy, with detailed results available in Appendix A.3.
@@ 5.1 Benchmark Quality Measurements
In this section, we delve into a detailed analysis of the benchmarks based on different measurements. Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) Vania et al. (2021) and Fisher information Lord (1980).
@@ 5.2 Benchmark Analysis
Based on the measurements, we present an aggregate rank for each benchmark based on the sum of individual ranks in Table 3, where a lower total score indicates better overall quality. We also present the distribution of item-level properties across 11 LLM benchmarks in Figure 3. Moreover, we further illuminates the relationship between item difficulty and discriminability in Figure 4. Based on these experiment results, we present our detailed analysis as follows.
@@
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data Zhuang et al. (2024).
@@
Feasibility estimates how well a question can be answered based on the information provided. Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear or defensible answer. As shown in Figure 3, datasets like TheoremQA and GPQA Diamond exhibit lower feasibility, likely due to complex scientific contexts or vague problem descriptions. In contrast, ARC-C and HellaSwag have high feasibility, suggesting their questions are well-formed and specific. Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities Rodriguez et al. (2021).
Finding 6: Item discriminability is weakened by excessively high or low item difficulty.
As shown in Table 3, benchmarks such as MATH and GSM8K achieve high discriminability, while GPQA Diamond and TheoremQA exhibit the lowest discriminability. This limited variability in model outcomes on these latter datasets may arise from too high item difficulty, thereby reducing their capacity for precise model differentiation.
Moreover, as shown in Figure 4, item discriminability tends to be lower when item difficulty reaches extreme levels—being either too easy or too hard. Consequently, datasets that feature a balanced distribution of difficulty often exhibit higher discriminability, enabling them to more effectively distinguish between models of varying capabilities.
@@ 6 Efficient Benchmarking with PSN-IRT
The pursuit of comprehensive LLM evaluation has often led to the creation of increasingly large and diverse benchmark datasets. However, the sheer volume of benchmark items does not inherently guarantee higher quality or more insightful evaluations, and can incur significant computational costs. In this section, we explore how PSN-IRT can facilitate more efficient benchmarking by guiding the strategic curation of smaller, yet highly informative, item sets that maintain or even enhance evaluation accuracy and reliability compared to larger, undifferentiated collections.
To empirically examine this principle, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involve Benchmark Agreement Testing (BAT) Perlitz et al. (2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena Chiang et al. (2024) and OpenCompass Arena Contributors (2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities.. Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess differentiation among stronger contenders.
As shown in Table 4, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking. For instance, a carefully selected subset composed of just 1000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048. This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
In contrast, other item selection metrics such as discriminability proves less effective for differentiating among strong models. While such items are typically adept at distinguishing between broad capability tiers (e.g., strong versus weak models), they often lack the granularity to capture nuanced performance differences at the higher end of the ability spectrum. This further verify the estimation reliability of our proposed PSN-IRT.
Metric w/ weak models w/o weak models

### v2

@@ Abstract
The evaluation of large language models (LLMs) via benchmarks is widespread, yet inconsistencies between different leaderboards and poor separability among top models raise concerns about their ability to accurately reflect authentic model capabilities. This paper provides a critical analysis of benchmark effectiveness, examining mainstream prominent LLM benchmarks using results from diverse models. We first propose Pseudo-Siamese Network for Item Response Theory (PSN-IRT), an enhanced Item Response Theory framework that incorporates a rich set of item parameters within an IRT-grounded architecture. PSN-IRT can be utilized for accurate and reliable estimations of item characteristics and model abilities. Based on PSN-IRT, we conduct extensive analysis on 11 LLM benchmarks comprising 41,871 items, revealing significant and varied shortcomings in their measurement quality. Furthermore, we demonstrate that leveraging PSN-IRT is able to construct smaller benchmarks while maintaining stronger alignment with human preference.
Codehttps://github.com/Joe-Hall-Lee/PSN-IRT
@@ 1 Introduction
As the scale and performance of large language models (LLMs) continue to grow, accurately measuring their capabilities has become increasingly important. Currently, the performance of LLMs is primarily evaluated through various benchmarks, which are comprehensive test suites consisting of carefully designed questions to assess model behavior across different tasks. However, in practice, existing benchmarks often exhibit significant limitations, prompting consideration of their effectiveness (McIntosh et al. 2024).
@@
As illustrated in Figure 2, on one hand, different benchmarks often produce inconsistent leaderboards (Perlitz et al. 2024). This issue persists even for benchmarks designed to evaluate similar underlying capabilities, leading to the same model achieving substantially varying rankings across evaluations. On the other hand, many benchmarks show weak separability among top models, limiting the ability to understand performance differences and further model refinement. Given these limitations, there is a growing need for a more systematic analysis of LLM benchmarks.
In this paper, we conduct a comprehensive analysis for various popular LLM benchmark datasets. Our experiment is based on Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004), a psychometric framework widely used in educational assessment, analyzes item properties such as difficulty by modeling item response against examinee ability [^3] In the paper, an item refers to a benchmark sample.. However, traditional IRT struggles with the complexity and scale of modern datasets, and its assumption of the normal distribution of abilities does not always hold.
@@
To validate the effectiveness of PSN-IRT, we conduct extensive experiments on evaluation results from 12 LLMs across 11 benchmarks. Our results show that PSN-IRT outperforms previous approaches in both parameter estimation accuracy and reliability. Based on PSN-IRT, we perform in-depth analyses of these benchmarks using key metrics, with main findings as follows:
@@
    We propose PSN-IRT, a benchmark analysis framework with superior estimation accuracy and reliability.
@@
    We use PSN-IRT for an in-depth analysis of mainstream benchmarks, and found that current LLM benchmarks present deficiencies in many aspects.
@@ 2.1 Related Work
Researchers have explored various methods to evaluate and analyze the quality of LLM benchmarks. Recent efforts have introduced quantitative metrics that operate at a holistic, dataset level. For example, Benchmark Agreement Testing (BAT) assesses benchmark reliability by comparing the consistency of model rankings across different leaderboards (Perlitz et al. 2024; White et al. 2025). Other novel metrics have also been proposed for more nuanced dataset-level analysis (Li et al. 2024). Delving deeper than macro-level comparisons, another significant line of work focuses on content-based analysis to ground evaluation in concrete capabilities. For instance, ADeLe (Zhou et al. 2025) leverages scalable cognitive rubrics and item demand features to build highly interpretable capability profiles for AI systems.
Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968; Baker and Kim 2004) offers a psychometric framework for data-driven, item-level analysis. Within NLP, IRT has been primarily utilized to diagnose fundamental item properties such as difficulty and discriminability (Vania et al. 2021; Byrd and Srivastava 2022). More recently, its principles have also been used to inspire methods for improving evaluation efficiency, for instance, through optimized item selection (Maia Polo et al. 2024) or adaptive testing paradigms (Zhuang et al. 2024). However, due to the complexity and scale of modern datasets and its assumption of normally distributed abilities, the application of IRT on LLM benchmarks is still underexplored (Ye et al. 2025).
@@ 2.2 Preliminary: Item Response Theory
Item Response Theory (IRT) (Lord, Novick, and Birnbaum 1968) is a psychometric framework widely used in educational and cognitive assessments to model the relationship between benchmark items and examinee abilities. Unlike Classical Test Theory (CTT) (DeVellis 2006), which relies on aggregate test scores, IRT analyzes individual item responses as a function of a latent ability $\theta$, enabling precise measurement of both item and examinee properties.
@@
More advanced IRT models introduce additional parameters to capture diverse item properties. However, traditional IRT often suffer from inaccurate parameter estimation and rely on assumptions like normal ability distributions, which may not align with real-world data (Tsutsumi, Kinoshita, and Ueno 2021).
@@ 3 Pseudo-Siamese Network for IRT
  • The discriminability $a$ reflects an item’s power to differentiate between models of varying abilities.
@@
4 Efficacy of PSN-IRT
4.1 Experimental Setup
Datasets.
@@ 4.1 Experimental Setup
To construct training data, we conduct evaluations using OpenCompass (Contributors 2023). The evaluation results are collected in the form of binary outcomes, indicating whether each model answered each item correctly. Then, we construct a unified outcome matrix by aggregating item-level results across all models and benchmarks. We split the interactions into training, validation, and test sets.
@@
ARC-C (Clark et al. 2018) 295 General EM Multiple Choice
BBH (Suzgun et al. 2023) 6,511 General EM Mixed
Chinese SimpleQA (He et al. 2024) 3,000 Knowledge LLM-as-a-Judge QA
GPQA Diamond (Rein et al. 2024) 198 Science EM Multiple Choice
GSM8K (Cobbe et al. 2021) 1,319 Math EM QA
HellaSwag (Zellers et al. 2019) 10,042 General EM Multiple Choice
HumanEval (Chen et al. 2021) 164 Code Pass@1 QA
MATH (Hendrycks et al. 2021b) 5,000 Math EM QA
MBPP (Austin et al. 2021) 500 Code Pass@1 QA
MMLU (Hendrycks et al. 2021a) 14,042 General EM Multiple Choice
TheoremQA (Chen et al. 2023) 800 Science EM QA
@@
Model Parameter Method ACC F1 AUC Kendall’s $\tau$
IRT 4PL MLE 0.7211 0.8034 0.7012 0.9697
MCMC 0.7070 0.7811 0.7278 0.9697
VI 0.7201 0.8015 0.6940 0.9091
VIBO 0.7188 0.8007 0.7055 0.9697
Deep-IRT 1PL Deep Learning 0.7974 0.8516 0.8519 0.9697
PSN-IRT 4PL Deep Learning 0.7998 0.8538 0.8485 1.0000
@@
We mainly focus on currently high-performing models, including 360GPT2-Pro [^4] https://ai.360.cn/, DeepSeek-V3 (DeepSeek-AI et al. 2025), Doubao-Pro [^5] https://www.volcengine.com/product/doubao, Gemini-1.5 (Team et al. 2024a), Hunyuan-Turbo [^6] https://hunyuan.tencent.com/, Moonshot-v1777https://platform.moonshot.cn/, Qwen-Plus (Qwen et al. 2025), and Yi-Lightning (Wake et al. 2025). However, testing only strong models might obscure differences in the discriminability of the test sets, as most items could be solved easily, thus limiting a comprehensive evaluation of test set quality (Mart´ınez-Plumed et al. 2019). To address this, we intentionally included several relatively weaker models, including Gemma-2B-it (Team et al. 2024b), Mistral-7B-Instruct-v0.1 (Jiang et al. 2023), Qwen2.5-3B-Instruct (Qwen et al. 2025), and Vicuna-7B-v1.3 (Chiang et al. 2023).
Baselines.
We mainly compare PSN-IRT with traditional IRT methods, namely IRT estimated based on traditional parameter estimation techniques, including Maximum Likelihood Estimation (MLE), Markov Chain Monte Carlo (MCMC) (Hastings 1970), Variational Inference (VI) (Jordan et al. 1999), and VIBO (Wu et al. 2020). We also compare with Deep-IRT (Tsutsumi, Kinoshita, and Ueno 2021), which is an extension of 1PL IRT that learns item-trait interactions via deep networks.
Metrics.
To assess whether the learned parameters meaningfully reflect corresponding properties, we use two quantitative metrics:
  • Prediction accuracy: Following standard practice in educational testing (Eignor 2013), we assess how well a method predicts whether a model answers an item correctly, reporting accuracy, F1 score, and ROC AUC.
  • Rank stability: To evaluate the reliability of model ranking induced by each method, we split the test set into two subsets, estimate model abilities separately, and compute Kendall’s $\tau$ between the resulting rankings.
4.2 Main Results
@@
Given our goal to comprehensively capture diverse item characteristics, we adopt the more expressive 4PL structure for all methods except Deep-IRT. Experiments also confirm that PSN-IRT performs best with 4PL, with detailed results provided in the Appendix B.1.
4.3 Ablation Study
To assess whether more complex architectures offer benefits over our straightforward design, we conduct an ablation study. Specifically, we compare PSN-IRT with two variants: one replaces one-hot inputs with semantic embeddings from Instructor (Su et al. 2023), and the other uses a GNN (Su et al. 2022) instead of the MLP backbone [^8] Please refer to Appendix A.2 for more details..
As shown in Table 3, both alternatives underperform our original design. We argue that semantic similarity between item texts does not necessarily reflect similarity in measurement characteristics. For the GNN variant, its neighborhood aggregation mechanism risks diluting the unique statistical signals of individual models and items, thereby hindering precise parameter estimation. These results suggest that in this estimation scenario, preserving the individuality of models and items is more effective than introducing architectural complexity.
Method ACC F1 AUC Kendall
PSN-IRT 0.7998 0.8538 0.8485 1.0000
   + Embedding 0.7808 0.8413 0.8310 0.9394
   + GNN 0.7928 0.8490 0.8197 0.7273
Table 3: Ablation study on PSN-IRT’s input representation and network architecture.
Furthermore, since PSN-IRT estimates parameters that are specific to the training data, requiring retraining to analyze new items, we also investigate the stability of these estimates when the item pool changes. To simulate this, We conduct an experiment by independently training separate PSN-IRT models on random subsets of the items. We then compute the Pearson correlation between the item parameters estimated from these subsets and those estimated from the full dataset.
As shown in Table 4, the results reveal a high degree of correlation across all parameters. This indicates that PSN-IRT learns highly consistent intrinsic properties for the items regardless of the specific data sample, demonstrating the stability of parameter estimation.
Data Difficulty Discrim. Guessing Feasibility
30% 0.9009 0.8186 0.8274 0.9025
50% 0.7437 0.7835 0.8776 0.9330
70% 0.9519 0.7414 0.8320 0.9442
Table 4: Pearson correlation coefficients of item parameters estimated on random subsets of the data versus those estimated on the full dataset. All correlations are statistically significant ($p<0.0001$).
@@ 5.1 Benchmark Quality Measurements
In this section, we delve into a detailed analysis of the benchmarks based on different measurements. Specifically, we leverage the 4 item parameters (difficulty, discriminability, guessing-rate and feasibility) estimated by PSN-IRT, along with two additional metrics for evaluating item quality: Local Efficiency Headroom (LEH) (Vania et al. 2021) and Fisher information (Lord 1980).
@@ 5.2 Benchmark Analysis
Based on the measurements, we provide an aggregate rank for each benchmark, calculated from the sum of individual ranks in Table 5, where a lower total score indicates better overall quality. Additionally, Figure 3 visualizes the distribution of item-level properties across the 11 LLM benchmarks. Building on these experimental results, our detailed analysis is presented as follows.
@@
The guessing-rate reflects the chance that a model can answer a question correctly without actual understanding, typically by exploiting format-based cues or shortcuts. As shown in Figure 3, datasets like ARC-C, HellaSwag, and MMLU generally present higher guessing-rates, possibly due to their multiple-choice formats. Notably, further inspection of the item guessing-rate parameter distributions, as depicted in Figure 3, reveals that a considerable number of items across several benchmarks exhibit unusually high guessing-rates. This observation regards the potential for such items to indicate data contamination, namely the content or answers for these items might have been present in the models’ pre-training data (Zhuang et al. 2024).
@@
Feasibility estimates how well a question can be answered based on the information provided. Low feasibility often arises from underspecified or overly broad questions, where even a strong model cannot determine a clear answer. As shown in Figure 3, datasets like TheoremQA and GPQA Diamond exhibit lower feasibility, likely due to complex scientific contexts or vague problem descriptions. In contrast, ARC-C and HellaSwag have high feasibility, suggesting their questions are well-formed and specific. Low feasibility more often reflects weaknesses in annotation quality or task design, and such items can distort evaluation by penalizing models for failing to resolve ambiguities (Rodriguez et al. 2021).
@@ 6 Efficient Benchmarking with PSN-IRT
While the pursuit of comprehensive LLM evaluation has led to increasingly large benchmarks, sheer item volume neither guarantees quality nor comes without significant computational cost. This section, therefore, explores how PSN-IRT can facilitate more efficient benchmarking. We investigate whether strategically curating smaller, highly informative item sets can yield model comparisons that maintain or even surpass the accuracy and reliability of larger, undifferentiated collections.
Experimental Setup.
To evaluate the effectiveness of different item selection strategies, we design an experiment to evaluate the effectiveness of different item selection strategies. Our methodology involve Benchmark Agreement Testing (BAT) (Perlitz et al. 2024). We first establish a reference model ranking by aggregating results from two public leaderboards: Chatbot Arena (Chiang et al. 2024) and OpenCompass Arena (Contributors 2023) [^9] These arenas are based on human evaluation and thus reflect human preference regarding overall model capabilities.. Subsequently, various subsets of benchmark items are chosen based on different criteria. Model rankings generated using these subsets are then compared against the reference arena ranking using Kendall’s $\tau$ to measure agreement. Alongside agreement, we also consider the variance of model scores on these subsets; higher variance generally indicates the subset’s capacity to differentiate more clearly between models. The evaluations are conducted using two distinct model groupings: one set comprising a mix of stronger and weaker models ("w/ weak models"), and another set from which weaker models are excluded ("w/o weak models"), to specifically assess separability among stronger contenders.
Results.
As shown in Table 6, selecting items based on Fisher information consistently produces model rankings with superior alignment to the reference arena ranking.For instance, a carefully selected subset composed of just 1,000 items chosen via the Fisher information criterion achieves a Kendall’s $\tau$ of up to 0.9048. This level of agreement markedly surpasses that achieved by rankings derived from using all available items or from similarly sized randomly selected subsets.
Method w/ weak models w/o weak models